PERISHER PTY LTD

Perisher Pty Ltd (Perisher) is a Ski equipment manufacturer that operates around Mt Hotham in Victoria. On 1 May 2019, Perisher provided Nikita (one of its employees) with a car as Nikita does a lot of travelling for work purposes. However, Nikita’s usage of the car is not restricted to work only. Perisher purchased the car on that date for $44,000 (including GST) plus $2,000 (including GST) dealer delivery charges. For the period of 1 May 2019 to 31 March 2020, Nikita travelled 12,000 kilometers in the car and incurred expenses of $770 on minor repairs that have been reimbursed by Perisher. The car was not used for 10 days when Nikita was interstate and was parked at the airport and for another five days when the car was scheduled for annual repairs. Calculate the Fringe Benefits Tax Liability for Perisher, please have a look at the matrix below on how to answer the question
Answer #1

Fringe benefit taxes are always paid by empoloyer if he provides extra benefit for one of his employees.

Car total value = 44000 + 2000 = $46000.

No. Of days the particular service is provided to the employee is from 1st may 2019 to 31 March 2020. = 304 days .

Fringe tax chargable at --30%

So at last the fringe tax payable is

30% of $46,000 × (304\365) - (expenses)$770

= 13800×[304/365] - 770 = $10, 723.6986

= $ 10,723.70 (approximately)
Answer #2

We can go through the answer step by step

The formula for the solution is,

Taxable value= (A*B*C)/D - E

Here,

A= Base value

B= Statutory percentage

C= Days for the available private uses

D= The number of days in FBT year

E= Employees contribution

And the basevalue means the orginal price + delivery charges

Thus, base value is =44000+2000= 46000

A flat statutory rate of 47% applies ending 31st march 2019

The car is treated as private use but it is not available at the employees premises, and the next time at workshop for the maintenances.

Thus at this both cases of parking at the airport and repair shop it can be considered as the car is available for the private uses.

Available for the private use is for, 10+5= 15 days

The FBT days is total 335 days

That is from 1st may 2019 to 31st march 2020

The employees contribution is 770

Thus the taxable value is,

= (A*B*C)/D - E

=(46000*47%*15)/335-770

=198.05